OpenStudy (anonymous):

Determine the points at which the graph of the following function has horizontal tangent. y = x^2/X^2-4

OpenStudy (anonymous):

take the derivative, set = 0 and solve.

OpenStudy (anonymous):

derivative is \[\frac{-8x}{(x^2-4)^2}\] so set \[-8x=0\] and get \[x=0\]

myininaya (myininaya):

\[f'(x)=\frac{2x(x^2-4)-x^2(2x)}{(x^2-4)^2}=\frac{2x(x^2-4-x^2)}{(x^2-4)^2}=\frac{2x(-4)}{(x^2-4)^2}\]

OpenStudy (anonymous):

my myininaya will write the steps

OpenStudy (anonymous):

oh right, before i can even finish

OpenStudy (anonymous):

so point is \[(0,0)\]